OpenStudy (anonymous):

Solve for x: 8x^2 + 64x = 0 x = 0, x = −8 x = 0, x = 8 x = −8, x = 0, x = 8 x = 8, x = 64

OpenStudy (anonymous):

@Haseeb96

OpenStudy (anonymous):

@tester97

OpenStudy (anonymous):

its a do u know how ?

OpenStudy (anonymous):

No can you please explain it?

OpenStudy (anonymous):

okay its very simple 8x^2 + 64x = 0 8x ( x + 8 ) =0 either 8x = o then x= 0 or x+8 = 0 then x = -8 x = {0 , -8}

OpenStudy (anonymous):

Thank you! Can you help me with a few more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Harry is trying to solve the equation y = 2x^2 − x − 6 using the quadratic formula. He has made an error in one of the steps below. Find the step where Harry went wrong. Step 1: x equals the negative of negative 1 plus or minus the square root of the quantity negative one squared minus 4 times 2 times negative six, end quantity, all over 2 times 2. Step 2: x equals the negative of negative 1 plus or minus the square root of negative one plus forty-eight all over two times 2. Step 3: x equals the negative of negative 1 plus or minus the square root of forty-seven all over two times 2. Step 4: x equals 1 plus or minus the square root of forty-seven all over 4. Step 1 Step 2 Step 3 Step 4

OpenStudy (anonymous):

@Hamoody1996

OpenStudy (anonymous):

2

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Solve for x: 25x^2 − 36 = 0 x = 6, 5 x = −6, −5 x = 36 over 25, x = − 36 over 25 x = 6 over 5, x = − 6 over 5

OpenStudy (anonymous):

x = 6 over 5, x = − 6 over 5

OpenStudy (anonymous):

How can you tell when a quadratic equation has two identical, rational solutions? When the radicand is negative When b in the quadratic formula is greater than the radicand When the radicand equals zero When the radicand is not a perfect square